Three particles with equal positive charges q are at the corners of an equilateral triangle of side a (a) at what point, if any,in the plane of the particles is the electric potential zero?(b) what is the electric potential at the position of one of the particles due to the other two particles in the triangle?
1
Expert's answer
2020-03-23T12:25:46-0400

a)

"V_{total}=k\frac{q}{r_1}+k\frac{q}{r_2}+k\frac{q}{r_3}=kq(r_1^{-1}+r_2^{-1}+r_3^{-1})>0"

There is not such point with zero potential.

b)


"V_{1}=k\frac{q}{a}+k\frac{q}{a}=2k\frac{q}{a}"